# CERN-TH-2024-177

### v1

@@ I Introduction
In this Letter, we discuss a novel approach to fit experimental scattering data using a UV complete ansatz that embodies the physical constraints of Analyticity, Crossing symmetry, and Unitarity. [^1] To the best of our knowledge, a fit ansatz satisfying all these constraints has never been constructed. We propose using non-perturbative S-matrix Bootstrap methods to construct a fit model that satisfies all those properties, with only a few tunable parameters. Those parameters are then fixed by minimizing the $\chi^{2}$ with the experimental data. We will call this strategy “Bootstrap Fit”.
@@
Pions are the lightest states of the QCD spectrum. A two-particle $\pi\pi$ state has $G$-parity $G=+1$, and parity $P=(-1)^{J}$, where $J$ is the spin of the particle. Only resonances with the same quantum numbers can be produced in a scattering experiment. In figure 1, the experimentally observed spectrum below $1.4$ GeV is represented with gray boxes, adding only the $\rho_{3}$ above this energy. [^2] To express the spectrum in dimensionless units we use $m_{\pi}=137.3$ MeV, the average between charged and neutral pions. The vertical size of the boxes sets the error on the mass, the horizontal size represents the width of the corresponding particle (the error on the width is not shown).
@@
The Numerical Bootstrap of pion scattering amplitudes was recently revived in Guerrieri et al. (2019) and later extended to the massless case in Guerrieri et al. (2021a). In Bose et al. (2020), the authors have refined the allowed region found in Guerrieri et al. (2019) using additional positivity constraints at low energy. By performing hypothesis testing using relative entropy, they have selected the region of low energy parameters that would match QCD low energy data and $\chi$PT. In Bose et al. (2021), they have also extracted the spectrum finding hints of emerging Regge trajectories. In He and Kruczenski (2023, 2024), the authors have used several low energy constraints below the two-particle threshold, and form factors constraints computed using perturbative QCD at energy above 2 GeV, to select a region in parameter space that nicely agrees with experimental phase shifts and low energy data. The Bootstrap of large-$N$ $\pi\pi$ amplitudes was kicked off in Albert and Rastelli (2022) and Fernandez et al. (2023), and generalized by including photons and matching with the chiral anomaly in Albert and Rastelli (2024). Finally, in Albert et al. (2024), including a minimal amount of spectrum assumptions, a Bootstrap candidate for the large-$N$ QCD amplitude was found.
@@
Our main result is the construction of a full-fledged analytic amplitude ${\cal T}_{ab}^{cd}(s,t,u)$ that fits the experimental data. From this amplitude, we predict the spectrum and low energy parameters. In figure 1, we present in red the Bootstrap estimate of the spectrum. Both the mass and the width of the resonances match the experimental measurements, except those denoted by an asterisk that are still affected by numerical systematics. The last column has the exotic quantum numbers $2^{+}(0^{++})$. There is no such state in the PDG Navas et al. (2024). It is a genuine prediction of our Bootstrap procedure. Before discussing in detail these results, we shall first describe the two steps of our strategy.
@@
In section II, we discuss the S-matrix Bootstrap step in constructing a candidate amplitude $\mathcal{A}^{\text{ansatz}}_{\Theta}(s|t,u)$ depending on a small set of free parameters $\Theta$ by solving a semi-definite optimization problem. Next, we construct and minimize the $\chi^{2}(\Theta)$ using a gradient-free Particle Swarm Optimization (PSO) algorithm, reviewed in section III. To evaluate the $\chi^{2}(\Theta)$ function we use the result of the Bootstrap problem $\mathcal{A}^{\text{ansatz}}_{\Theta}(s|t,u)$, and the experimental and lattice data available for the $S_{0}$, $S_{2}$, $P$, and $D_{2}$ waves, as explained in section IV, where we present the result of the Bootstrap Fit. In section V we test the amplitude obtained in different regimes by comparing it with the available experimental measurements, finding always a good qualitative agreement. Finally, in section VI we comment on possible improvements and future directions.
@@ II.2 The unitarity constraints
As for the functions $\eta_{\ell}^{(I)}(s)$, we use the phenomenological parametrization discussed in Garcia-Martin et al. (2011). These functions should be regarded as an additional phenomenological input. [^7] It would be interesting to consider the general mixed system of pions and kaons, and obtain the inelasticity as a result of the procedure. See Homrich et al. (2019); Bercini et al. (2020); Guerrieri et al. (2020) for S-matrix Bootstrap works on mixed-amplitudes.
@@ II.3 Soft theorems and Spectrum Assumptions
Soft theorems are the consequence of spontaneously broken chiral symmetry. If pions were massless, the amplitude would vanish $\mathcal{A}\to 0$ when any of the momenta of the particles become soft. As the pions are massive, this equation is corrected by quantum effects and is no longer exact Adler (1965). Nevertheless, the existence of low energy zeroes in the partial waves is a prediction of Chiral Perturbation Theory ($\chi$PT) Weinberg (1979). For this reason, we impose that $t_{0}^{(0)}(z_{0})=0$, and $t_{0}^{(2)}(z_{2})=0$ for some $0\leq z_{0},z_{2}\leq 4$. We refer to these constraints as “chiral zeroes” conditions. [^8] At tree-level in $\chi$PT: $z_{0}=1/2$, and $z_{2}=2$.
The physical spectrum is encoded into the position of pole singularities in the second sheet of the $2\to 2$ scattering amplitude. Using the elastic unitarity condition, it is possible to relate those resonance poles to zeros of the $S$-matrix (not the amplitude!) in the first sheet. [^9] When resonances have a mass higher than multi-particle thresholds the analytic continuation becomes subtle. While resonance poles imply zeros in the first sheet, in general, zeros do not have to be the signature of a real resonance state. We impose several conditions of the form $S_{\ell}^{(I)}(s_{R})=0$, that we call “resonance zeros”, where $s_{R}$ is the complex mass squared of the particle, and $(\ell,I)$ its quantum numbers.
At this point, it is important to stress that we are agnostic about the values of the chiral and resonance zeros. In our construction, we only assume their existence with the corresponding quantum numbers. Their numerical values will be a prediction of the fit.
For the spectrum, we only assume the existence of a part of it. It turns out that to correctly reproduce the experimental data, we will need to consider at least four resonance zero constraints, one for the $\rho(770)$ in the $P$ wave, the two $f_{0}$’s in the $S_{0}$, [^10] We call the resonances $f_{0},f_{0}^{\prime},f_{0}^{\prime\prime},\dots$ to differentiate them. We will use similar notation in other channels. In the $S_{0}$ channel, we do not include the $\sigma$ in this list. and one $f_{2}$ in the $S_{2}$ wave. We will then discover the full spectrum by searching for additional dynamically generated zeros in the partial waves.
@@ II.4 The target functional
The final step in building our fit model is choosing the Bootstrap target functional. There are in principle infinitely many options, but not all are equivalent. For our purpose, we have found it convenient to follow the objective studied in Guerrieri et al. (2019). [^11] It would be interesting to understand the relation between this objective and the regions selected in Bose et al. (2020, 2021), and He and Kruczenski (2023, 2024) using alternative principles. To explain the why, let us consider the most general ansatz (3), and fix the $S_{0}$ scattering length $a_{0}^{(0)}$ [^12] Scattering lengths are defined from the threshold expansion of the partial amplitudes $\mathop{\mathrm{Re}}t^{(I)}_{\ell}(s)=2k^{2\ell}(a_{\ell}^{(I)}+k^{2}b_{\ell}^{(I)}+\dots)$, where $k^{2}=s/4-1$ is the center of mass momentum. to a certain value, as well as for the chiral zeros $z_{0}$ and $z_{2}$. If we then look at the allowed region of the $\{a_{1}^{(1)},a_{0}^{(2)}\}$ scattering lengths we generically find a kink as in figure 2. Tuning the values of the $\{a_{0}^{(0)},z_{0},z_{2}\}$ parameters, it is possible to obtain a kink close to the experimental values of the scattering lengths. When that happens, the extremal amplitude at the kink closely resembles the pion amplitude, but not exactly. [^13] The values of scattering lengths, and subleading threshold coefficients match the experimental values, and the corresponding amplitude contains both the $\sigma$, and the $\rho$ resonance states. However, although the $\sigma$ position is compatible with the data, the width of the $\rho$ is larger than the experimental one. We have tried to fine tune the triplet $\{a_{0}^{(0)},z_{0},z_{2}\}$ in order to match the experimental values of $\sigma$ and $\rho$, but without success. We conjecture that by fine-tuning additional low energy constants it would be possible to dynamically generate the $\rho$ in the correct position. For this reason, we expect that imposing a spectrum assumption might be enough to obtain a good fit model for the pion amplitude.
Figure 2: The kink from Guerrieri et al. (2019) in the scattering length space $\{a_{0}^{(0)},a_{1}^{(1)}\}$. The allowed region is in blue. The point with error bars is the experimental determination, see also Table 2. The blue dot is the position of the kink as a result of our Bootstrap fit algorithm.
The Bootstrap problem designed to construct the fit model of the pion amplitude can be summarized as follows. {mdframed}[frametitle=Fit Ansatz,frametitlealignment=,backgroundcolor=blue!6, leftmargin=0cm, rightmargin=0cm, topline=false,bottomline=false, leftline=false, rightline=false]
@@
The objective is the “kink functional”
@@
which depends on the angle $\theta$ that typically is somewhere between $\pi<\theta<3\pi/2$. We let $\theta$ free, and add it to the set of fit parameters. [^14] We could consider a different objective $\text{Obj}^{\prime}(\theta,\phi)=a_{0}^{(0)}\cos\phi\sin\theta+a_{0}^{(2)}\sin\phi\sin\theta+a_{1}^{(1)}\cos\theta$, and replace the $a_{0}^{(0)}$ parameter with a new angle $\phi$. This would lead to an equivalent formulation of the Bootstrap ansatz (9). From this view, it is evident we are moving along the two-dimensional boundary of the three-dimensional parameter space of the scattering lengths. This can be generalized by adding an arbitrary number of low energy observables, which amounts to scanning over a larger space of amplitudes.
The quantities collectively denoted by $\Theta$ are the parameters of the fit and the input for the optimization problem (9). We solve this problem using the standard SDPB solver Simmons-Duffin (2015); Landry and Simmons-Duffin (2019). The corresponding extremal amplitude depends on the choice of $\Theta$. Next, we will explain how to optimize on $\Theta$. For this problem, the size of the parameter set is $|\Theta|=12$.
@@ III Gradient free optimization and particle swarm
In this Letter, we explore an algorithm especially suited to this class of problems, the Particle Swarm Optimization algorithm (PSO) Shi and Eberhart (1998). [^15] We are grateful to Balt van Rees for pointing out this method. The PSO is a standard algorithm designed for solving non-convex problems, with a wide range of applications – see Clerc (2006) for an introduction. The PSO is also gradient-free and does not require the computation of derivatives of the Bootstrap solution in $\Theta$ [^16] As explained in Reehorst et al. (2021), it is possible to efficiently compute the gradient of a Bootstrap problem solved with the interior point method as the one implemented in SDPB.
The general idea is simple, but the implementation details may depend on the problem. We start with $n_{p}$ particles at random positions $\Theta^{(i)}_{0}$ to which we assign random velocities $v^{(i)}_{0}$. Positions are drawn from a distribution peaked in the region where we believe the fit parameters live. Velocities are of the same order as the size of this region.
@@
We denote with $X_{n}^{(i)}$ the position with the best $\chi^{2}$ found by the $i$-th particle along its trajectory up to the step $n$ and with $Y_{n}$ the position with the best $\chi^{2}$ found by the entire swarm. $r_{1,2}$ are random numbers uniformly distributed between $[0,1]$. To construct each step, we first evaluate the $\chi^{2}(\Theta_{n}^{(i)})$ for each particle, then we use this information to update the $X_{n}^{(i)}$, and $Y_{n}$.
Three parameters control the algorithm performance: the inertia $\omega$, the cognitive coefficient $c_{1}$, and the cooperation coefficient $c_{2}$. The convergence property of the algorithm depends on the choice of these three coefficients. [^17] In other versions of the algorithm, those parameters can be promoted to be dynamical. The details of our implementation of the PSO algorithm, which contains an additional adaptive velocity prescription as in Xu (2013), can be found in Appendix B.
The algorithm runs for $N_{\text{iter}}$ steps. The best position found by the swarm at the last step $Y_{N_{\text{iter}}}$ is our estimate for the best fit parameters which gives the minimum value of $\chi^{2}(\Theta)$.

### v2

@@ I Introduction
In this Letter, we discuss a novel approach to fit experimental scattering data using a UV complete ansatz that embodies the physical constraints of Analyticity, Crossing symmetry, and Unitarity. [^1] To the best of our knowledge, a fit ansatz satisfying all these constraints has never been constructed. The closest representation satisfying some of these properties is in terms of Roy equations Roy (1971); Ananthanarayan et al. (2001). We propose using non-perturbative S-matrix Bootstrap methods to construct a fit model that satisfies all those properties, with only a few tunable parameters. Those parameters are then fixed by minimizing the $\chi^{2}$ with the experimental data. We will call this strategy “Bootstrap Fit”.
@@
Pions are the lightest states of the QCD spectrum. A two-particle $\pi\pi$ state has $G$-parity $G=+1$, and parity $P=(-1)^{J}$, where $J$ is the spin of the particle. Only resonances with the same quantum numbers can be produced in a scattering experiment. In figure 1, the experimentally observed spectrum below $1.4$ GeV is represented with gray boxes, adding only the $\rho_{3}$ above this energy. [^2] To express the spectrum in dimensionless units we use $m_{\pi}=137.3$ MeV, the average between charged and neutral pions.
@@
The Numerical Bootstrap of pion scattering amplitudes was recently revived in Guerrieri et al. (2019) and extended to the massless case in Guerrieri et al. (2021a). In Bose et al. (2020), the authors refined the allowed region found in Guerrieri et al. (2019) using additional positivity constraints at low energy. By performing hypothesis testing using relative entropy, they selected the region of low energy parameters that matches QCD low energy data and $\chi$PT. In Bose et al. (2021), they have also extracted the spectrum finding hints of emerging Regge trajectories. In He and Kruczenski (2023, 2024), the authors used several low energy constraints below the two-particle threshold, and form factors constraints computed in perturbative QCD at energy above 2 GeV, to select a region in parameter space that nicely agrees with experimental phase shifts and low energy data. The Bootstrap of large-$N$ $\pi\pi$ amplitudes was kicked off in Albert and Rastelli (2022) and Fernandez et al. (2023), and generalized by including photons and matching with the chiral anomaly in Albert and Rastelli (2024). Finally, including a minimal amount of spectrum assumptions, a Bootstrap candidate for the large-$N$ QCD amplitude was found in Albert et al. (2024).
@@
Our main result is the construction of a full-fledged analytic amplitude ${\cal T}_{ab}^{cd}(s,t,u)$ that fits the experimental data. From this amplitude, we predict the spectrum and the low energy parameters. In figure 1, we present in red the Bootstrap estimate of the spectrum. Both the mass and the width of the resonances match the experimental measurements, except for those denoted by an asterisk still affected by numerical systematics. The last column has the exotic quantum numbers $2^{+}(0^{++})$. There is no such state in the PDG Navas et al. (2024). Because of the $I=2$ quantum number, this particle is a genuine Tetraquark state Jaffe (1977); ’t Hooft et al. (2008) with a mass of around 2 GeV and a width of 600 MeV. As a prelude to our results, in figure 2 we show the predicted phase shift in the $S_{2}$ wave. At first glance, it would be difficult to anticipate from the experimental data that the phase shift would reverse direction at higher energies. This behavior is a genuine prediction of our Bootstrap procedure which incorporates full crossing and unitarity of the scattering amplitude. Before discussing in detail the physics of this amplitude we shall describe the steps of our strategy.
Figure 2: High energy $S_{2}$ partial wave phase shift profile. In blue the best fit, in light blue all amplitudes with suboptimal $\chi^{2}$ (see section IV). We see a jump in the phase around $s\approx 200$ that signals the presence of a resonance. In the inset, the mass parameters are extracted from $m=\sqrt{s^{*}}$, where $s^{*}$ is the position of the zero $S_{0}^{(2)}(s^{*})=0$ in the complex $s$-plane, and in units of $m_{\pi}=1$.
@@
In section II, we discuss the S-matrix Bootstrap step in constructing a candidate amplitude $\mathcal{A}^{\text{ansatz}}_{\Theta}(s|t,u)$ depending on a small set of free parameters $\Theta$ by solving a semi-definite optimization problem. In section III, we describe the $\chi^{2}(\Theta)$ minimization procedure using a gradient-free Particle Swarm Optimization (PSO) algorithm. To evaluate the $\chi^{2}(\Theta)$ function we use the result of the Bootstrap problem $\mathcal{A}^{\text{ansatz}}_{\Theta}(s|t,u)$, and the experimental and lattice data available for the $S_{0}$, $S_{2}$, $P$, and $D_{0}$ waves, as explained in section IV, where we present the result of the Bootstrap Fit. In section V we test the amplitude obtained in different regimes by comparing it with Chiral Perturbation Theory ($\chi$PT) and the experimental measurements not used in the fitting procedure, always finding a good qualitative agreement. Finally, in section VI we comment on possible improvements and future directions.
@@ II.2 The unitarity constraints
As for the functions $\eta_{\ell}^{(I)}(s)$, we use the mean value of the phenomenological parametrization for the “big-dip” scenario discussed in Garcia-Martin et al. (2011). [^7] For simplicity, we do not take into account the error on this parametrization. These functions should be regarded as an additional phenomenological input. [^8] It would be interesting to consider the general mixed system of pions and kaons, and obtain the inelasticity as a result of the procedure. See Homrich et al. (2019); Bercini et al. (2020); Guerrieri et al. (2020) for S-matrix Bootstrap works on mixed-amplitudes.
@@ II.3 Soft theorems and Spectrum Assumptions
Soft theorems are the consequence of spontaneously broken chiral symmetry. If pions were massless, the amplitude would vanish $\mathcal{A}\to 0$ when any of the momenta of the particles become soft. As the pions are massive, this behavior is corrected by quantum effects and is no longer exact Adler (1965). Nevertheless, the existence of low energy zeroes in the partial waves is a prediction of $\chi$PT Weinberg (1979). For this reason, we impose that $t_{0}^{(0)}(z_{0})=0$, and $t_{0}^{(2)}(z_{2})=0$ for some $0\leq z_{0},z_{2}\leq 4$. We refer to these constraints as chiral zeroes conditions. [^9] At tree-level in $\chi$PT: $z_{0}=1/2$, and $z_{2}=2$.
The physical spectrum is encoded into the position of pole singularities in the second sheet of the $2\to 2$ scattering amplitude. Using the elastic unitarity condition, it is possible to relate those resonance poles to zeros of the $S$-matrix (not the amplitude!) in the first sheet. We impose several conditions of the form $S_{\ell}^{(I)}(s_{R})=0$, that we call resonance zeros, where $s_{R}$ is the complex mass squared of the particle, and $(\ell,I)$ its quantum numbers.
At this point, it is important to stress that we are agnostic about the values of both the chiral and resonance zeros. In our construction, we only assume their existence with the corresponding quantum numbers. Their numerical values will be a prediction of the fit.
For the spectrum, we only assume the existence of a part of it. It turns out that to correctly reproduce the experimental data, we will need to consider at least four resonance zero constraints, one for the $\rho(770)$ in the $P$ wave, the two $f_{0}$’s in the $S_{0}$, [^10] We call the resonances $f_{0},f_{0}^{\prime},f_{0}^{\prime\prime},\dots$ to differentiate them. We will use similar notation in other channels. In the $S_{0}$ channel, we do not include the $\sigma$ in this list. and one $f_{2}$ in the $S_{2}$ wave. We will then discover the spectrum by searching for additional dynamically generated zeros in the partial waves.
@@ II.4 The target functional
The final step in constructing our fit model is selecting the Bootstrap target functional. While there are, in principle, infinitely many possible choices, not all are equivalent. For our purposes, we found it convenient to adopt the approach developed in Guerrieri et al. (2019). To illustrate why, consider the most general ansatz (3). Assigning specific values to the $S_{0}$ scattering length [^11] Scattering lengths are defined from the threshold expansion of the partial amplitudes $\mathop{\mathrm{Re}}t^{(I)}_{\ell}(s)=2k^{2\ell}(a_{\ell}^{(I)}+k^{2}b_{\ell}^{(I)}+\dots)$, where $k^{2}=s/4-1$ is the center of mass momentum. $a_{0}^{(0)}$ and the chiral zeros $z_{0}$ and $z_{2}$, we find that the allowed region for the $\{a_{1}^{(1)},a_{0}^{(2)}\}$ scattering lengths exhibits a kink, as shown in figure 3. Changing the values of the $\{a_{0}^{(0)},z_{0},z_{2}\}$ parameters, it is possible to move this kink closer to the experimentally measured scattering lengths (the red point in figure 3). When this occurs, the extremal amplitude at the kink closely resembles the pion amplitude.
In this case, the scattering lengths and subleading threshold coefficients align well with experimental data, and the corresponding amplitude contains both the $\sigma$, and the $\rho$ resonances. However, although the $\sigma$ position agrees with the data, the $\rho$ width is larger than observed experimentally. [^12] We tried to fine-tune the triplet $\{a_{0}^{(0)},z_{0},z_{2}\}$ to match the experimental values of $\sigma$ and $\rho$, but without success. We conjecture that by fine-tuning additional low energy constants it would be possible to generate the $\rho$ in the correct position dynamically. This suggests that by adding the spectral assumption, it may be possible to develop a robust fit ansatz for the pion amplitude.
Figure 3: The kink from Guerrieri et al. (2019) in the scattering length space $\{a_{0}^{(0)},a_{1}^{(1)}\}$. The allowed region is in blue. The point with error bars is the experimental determination, see also Table 2. The blue dot is the position of the kink as a result of our Bootstrap fit algorithm. On the left, we represent the target functional for $\theta=25\pi/18$.
The Bootstrap problem formulated to construct the fit model for the pion amplitude can be summarized as follows. {mdframed}[frametitle=Fit Ansatz,frametitlealignment=,backgroundcolor=blue!6, leftmargin=0cm, rightmargin=0cm, topline=false,bottomline=false, leftline=false, rightline=false]
@@
The objective we maximize is given by the linear combination
@@
which depends on the angle $\theta$. This objective is a normal functional and is effective at selecting “kinks” Córdova et al. (2020). To understand this, note that this objective maximizes the amplitude in the $\{a_{0}^{(0)},a_{1}^{(1)}\}$ space at a point where the tangent vector is orthogonal to $(a_{1}^{(1)}\cos\theta,a_{0}^{(2)}\sin\theta)$. At a kink, where multiple tangents exist, many normal functionals naturally converge to the kink point. For $\pi<\theta<3\pi/2$, this functional span the boundary of the allowed blue region in figure 3. In the left inset, we show a typical choice of angle $\theta$ that will select the kink in the $\{a_{0}^{(0)},a_{1}^{(1)}\}$ space. [^13] We could consider a different objective $\text{Obj}^{\prime}(\theta,\phi)=a_{0}^{(0)}\cos\phi\sin\theta+a_{0}^{(2)}\sin\phi\sin\theta+a_{1}^{(1)}\cos\theta$, and replace the $a_{0}^{(0)}$ parameter in (9) with a new angle $\phi$. This would lead to an equivalent formulation of the Bootstrap ansatz (9). From this view, it is evident we are moving along the two-dimensional boundary of the three-dimensional parameter space of the scattering lengths. This can be generalized by adding an arbitrary number of low energy observables, which amounts to scanning over a larger space of amplitudes.
The quantities collectively denoted by $\Theta$ are the parameters of the fit and the input for the optimization problem (9). In this problem, four parameters $\theta,a_{0}^{(0)},z_{0},z_{2}$ are real, while the mass square are complex. Their real and imaginary parts correspond respectively to the physical mass and width of the resonance. In total, the size of the parameter set in $|\Theta|=12$. We solve this problem using the standard SDPB solver Simmons-Duffin (2015); Landry and Simmons-Duffin (2019). The corresponding extremal amplitude depends on the choice of $\Theta$. Next, we will explain how to optimize on $\Theta$.
@@ III Gradient free optimization and particle swarm
In this Letter, we explore an algorithm especially suited to this class of problems, the Particle Swarm Optimization algorithm (PSO) Shi and Eberhart (1998). [^14] We are grateful to Balt van Rees for pointing out this method. The PSO is a standard algorithm designed for solving non-convex problems, with a wide range of applications—see Clerc (2006) for an introduction. The PSO is also gradient-free and does not require the computation of derivatives of the Bootstrap solution in $\Theta$ [^15] As explained in Reehorst et al. (2021), it is possible to efficiently compute the gradient of a Bootstrap problem solved with the interior point method as the one implemented in SDPB.
We start with $n_{p}$ particles at random positions $\Theta^{(i)}_{0}$ to which we assign random velocities $v^{(i)}_{0}$.
@@
The velocity of the $i$-th particle at step $n+1$ is thus a linear combination of its previous velocity, the distance to its position with the lowest $\chi^{2}$, denoted by $X_{n}^{(i)}$ and the distance with the overall best position among all particles $Y_{n}$. At each step, we first compute the $\chi^{2}(\Theta_{n}^{(i)})$ for each particle, then evaluate the various parameters to perform the step.
Three parameters control the algorithm performance: the inertia $\omega$, the cognitive coefficient $c_{1}$, and the cooperation coefficient $c_{2}$. [^16] The $r_{1,2}$ are random numbers uniformly distributed between $[0,1]$, and are drawn at each step. The convergence property of the algorithm depends on the choice of these three coefficients. [^17] In other versions of the algorithm, those parameters can be promoted to be dynamical. The details of our implementation of the PSO algorithm, which contains an additional adaptive velocity prescription as in Xu (2013), can be found in Appendix B.